OpenStudy (anonymous):

WHAT IS THE WEIGHT OF 1000 ATOMS OF NITROGEN

sam (.sam.):

Are you curious, by the way you can do that by just some fractions, we'll take diatomic Nitrogen here \[100atoms \times \frac{1molN_2}{6\times 10^{23}atoms} \times \frac{14(2)gN_2}{1molN_2}=?\] Work that out you will get grams of \(N_2\)